Two parallel lines are cut by a transversal. Angle 1 measures (4x + 28)°, and the angle adjacent to the alternate exterior angle with angle 1 measures (14x + 8)°. What is the value of x?
Answer by Cromlix(4381)

Hi there,
By your description of the
problem, it would appear
that if the angle adjacent to
the alternate angle with angle 1
measuring(14x + 8) is added to
the alternate angle with angle 1
they will = 180 degrees
(Both on straight line)
(4x + 28) + (14x + 8) = 180
Remove brackets
4x + 28 + 14x + 8 = 180
Collect like terms:
4x + 14x = 180 - 28 - 8
18x = 144
x = 8
Hope this helps :-)